I would like to see if I could get assistance to show me how to work this problem.
The probability of winning $20 in a particular lottery is 0.01. What is the probablity that you will get at least one $20 winner if you buy 90 tickets.
Thanks in advance!
Jennifer

Answer by stanbon(75887)

The probability of winning $20 in a particular lottery is 0.01. What is the probablity that you will get at least one $20 winner if you buy 90 tickets.
----
Binomial Problem with n = 90 and p = 0.01
----
P(at least one winner) = 1 - P(no winners)
---
P(at least one) = 1 - (0.99)^90 = 1- 0.4047 = 0.5953
==============
Cheers,
Stan H.